# STAvatar: Soft Binding and Temporal Density Control for Monocular 3D Head Avatars Reconstruction

### v2

@@ Abstract
Reconstructing high-fidelity and animatable 3D head avatars from monocular videos remains a challenging yet essential task. Existing methods based on 3D Gaussian Splatting typically bind Gaussians to mesh triangles and model deformations solely via Linear Blend Skinning, which results in rigid motion and limited expressiveness. Moreover, they lack specialized strategies to handle frequently occluded regions (e.g., mouth interiors, eyelids). To address these limitations, we propose STAvatar, which consists of two key components: (1) a UV-Adaptive Soft Binding framework that leverages both image-based and geometric priors to learn per-Gaussian feature offsets within the UV space. This UV representation supports dynamic resampling, ensuring full compatibility with Adaptive Density Control (ADC) and enhanced adaptability to shape and textural variations. (2) a Temporal ADC strategy, which first clusters structurally similar frames to facilitate more targeted computation of the densification criterion. It further introduces a novel fused perceptual error as clone criterion to jointly capture geometric and textural discrepancies, encouraging densification in regions requiring finer details. Extensive experiments on four benchmark datasets demonstrate that STAvatar achieves state-of-the-art reconstruction performance, especially in capturing fine-grained details and reconstructing frequently occluded regions. The code will be publicly available.
@@ 1 Introduction
Reconstructing animatable and photo-realistic 3D head avatars has been a longstanding problem in computer vision and graphics [egger20203d, saito2024relightable, wang20243d], driven by growing demands in AR/VR, telepresence, digital humans and interactive media. To achieve high-quality avatar reconstruction, conventional methods typically rely on expensive multi-camera systems [guo2019relightables, yang2023towards] which require complex manual setups and are impractical for consumer-level deployment. In contrast, monocular-based approaches [tang2025gaf, zheng2022avatar, grassal2022neural] aim to produce high-fidelity results using a single consumer-grade camera, enabling broader accessibility in immersive and personalized virtual environments.
@@
Recent years have witnessed the emergence of 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] as a powerful volumetric representation, delivering both high-fidelity rendering and real-time performance in neural scene reconstruction. Building on its success in static scenes, many works have explored its application to animatable 3D head avatar reconstruction. These methods typically follow a two-stage pipeline: a parametric head-tracking algorithm [qian2024vhap, zielonka2022towards] is first employed to estimate the facial pose, expression, and shape for each frame; subsequently, Gaussian primitives are bound to mesh triangles and driven by a deformation field, which is typically implemented using Linear Blend Skinning (LBS) [qian2023gaussianavatars]. During training, each Gaussian’s parameters are optimized via differentiable rendering to minimize photometric loss. Despite achieving enhanced visual fidelity and animation realism, existing Gaussian-based approaches still face several open challenges. To address these limitations, we propose STAvatar, as shown in Fig. 1, and provide a detailed explanation below.
@@
First, existing animatable reconstruction methods often suffer from rigid connection between Gaussian primitives and mesh triangles, resulting in poor handling of non‑rigid and fine‑scale deformations. Fig. 2(a) illustrates this simple hard binding framework [zhang2025fate, SplattingAvatar:CVPR2024] in which each Gaussian deforms solely according to the LBS of its parent triangle. Although this enables basic animation control, it fails to model fine-grained deformations, as the Gaussians remain relatively static within the triangle’s local coordinate frame. This limitation stems from the gap between mesh and Gaussian representations. While meshes inherently encode topology and support native skinning for deformation, Gaussians are structureless and rely on externally defined deformation fields to capture complex motions and fine details such as facial wrinkles. Some methods [chen2024monogaussianavatar, xiang2024flashavatar, giebenhain2024npga] attempt to enhance each Gaussian’s deformation using a fixed-dimension multilayer perceptron (MLP). However, this design requires predefining a fixed number of Gaussians, which is suboptimal since different identities often demand varying Gaussian densities. Besides, it neglects spatial context and may be incompatible with the Adaptive Density Control (ADC) inherent to 3DGS. To enhance deformation ability while preserving the flexibility of ADC, STAvatar proposes a UV-Adaptive Soft Binding framework, which integrates the LBS for coarse deformation and a dual-branch network for fine-detail recovery. By generating a feature offset map in UV space, STAvatar leverages spatial context to model the Gaussian deformation field and enables arbitrary-resolution sampling to support ADC.
@@
Second, the vanilla ADC in 3DGS is tailored for static scene reconstruction and fails to handle transiently visible or frequently occluded regions in dynamic avatar reconstruction. As illustrated in Fig. 2(b), certain regions (e.g., mouth interiors) are only visible in a subset of frames and contribute minimally to rendering when occluded. This results in low average criterion-such as the positional gradient used in 3DGS-which impedes sufficient densification. In addition, as shown in Fig. 2(c), the positional gradient in 3DGS only captures geometric discrepancies while neglecting texture details [rota2024revising], which leads to the omission of regions with high texture errors and the failure to add Gaussians to these regions. To address these limitations of ADC, STAvatar propose a Temporal Adaptive Density Control strategy incorporating FLAME-Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE-AP). As shown in Fig. 2(b), FTC partitions frames into structure‑coherent clusters to encourage frequently occluded regions remain visible within a cluster, facilitating targeted refinement. FPE‑AP replaces positional gradients by directly estimating each Gaussian’s fused perceptual error, which jointly considers both geometric and textural discrepancies. Besides, FPE-AP selects Gaussians exhibiting the highest instantaneous errors across iterations to better capture regions with large peak errors.
@@ 2.1 Gaussian Head Reconstruction
3DGS [kerbl3Dgaussians] has recently emerged as a flexible and efficient approach for 3D scene representation. By employing a large number of anisotropic Gaussian primitives—each parameterized by geometry and appearance attributes—it achieves high-quality reconstructions, making it particularly well-suited for detailed human head modeling. To enable avatar driving, typical methods [qian2023gaussianavatars, liu2025lucas, tang2025gaf, saunders2025gasp, SplattingAvatar:CVPR2024, zhang2025fate] adopt a hard-binding framework that associates Gaussians with mesh triangles to transfer mesh motion via LBS, which, however, fails to bridge the inherent gap between mesh and Gaussian representations, leading to suboptimal performance in handling complex motions and expressions. Alternatively, other approaches [xiang2024flashavatar, chen2024monogaussianavatar, xu2024gaussian, liao2023hhavatar] leverage MLP to predict per-Gaussian attribute offsets, thereby achieving a more flexible soft-binding framework. However, these methods often neglect ADC by using a fixed number of Gaussians across different identities, which limits expressiveness and reconstruction quality. In addition, several studies [ma20243d, li2025rgbavatar, zielonka2025gaussian] propose to represent avatars with a base model and a set of expression blendshapes. Although these methods can achieve high training efficiency, they often suffer from low fidelity or produce blurred results under large pose variations, since the learned blendshapes may not adequately capture diverse pose-dependent deformations.
@@ 2.2 Adaptive Density Control of 3DGS
ADC strategy is a key module in 3DGS which dynamically clones, splits and prunes Gaussians to balance detail preservation and representation compactness. The vanilla ADC selects primitives for densification based on average positional gradients and Gaussian scales. However, this heuristic often results in underfitting and blurred details, as it neglects texture-related cues and only focus average criterion. To overcome these limitations, some recent works [rota2024revising, zhang2024pixel, kim2024color, mallick2024taming, zhou2025perceptual] revise the densification criterion to incorporate richer information, thereby achieving finer details in underrepresented regions. Besides, optimization-theoretic approaches have been proposed to establish more principled rules for cloning and splitting: SteepGS [wang2025steepest] leverages steepest-descent analysis to trigger Gaussian splits, whereas 3DGS-MCMC [kheradmand20243d] formulates density control as a Markov Chain Monte Carlo (MCMC) sampling process with $L_{1}$-regularized pruning. More recently, Efficient Density Control [deng2024efficient] introduces long-axis splitting operations and recovery-aware pruning strategies to reduce redundant Gaussians and accelerate convergence, while Gradient-Direction-Aware ADC [zheng2025gradient] incorporates gradient coherence and angular alignment to more effectively determine which Gaussians should be cloned or split. However, as these methods are primarily tailored for static scene reconstruction, they tend to under-fit transiently visible regions and fail to jointly consider both geometric and textural discrepancies.
@@ 3 Method
We first introduce the general Gaussian avatar reconstruction pipeline in Sec. 3.1. Next, we present a UV-Adaptive Soft Binding framework in Sec. 3.2. The Temporal Adaptive Density Control strategy, which integrates FLAME‑Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE‑AP), is discussed in Sec. $3.$ Finally, the detailed optimization process is described in Sec. 3.4.
@@ 3.1 Preliminaries
Initialization. Following [qian2023gaussianavatars], we initialize Gaussian primitives $g_{i}$ based on the tracked FLAME model. Each $g_{i}$ is bound to a mesh triangle and parameterized by a center position $\bm{\mu}$, a scale matrix $\bm{S}$, and a rotation matrix $\bm{R}$ defined on the canonical FLAME mesh [li2017learning]. The 3D Gaussian density is defined as:
@@ 3.3 Temporal Adaptive Density Control
To estimate each Gaussian’s 2D projection for fused perceptual error calculation, we record its screen-space center $(x_{i},y_{i})$, total number of covered pixels $C_{i}$, and accumulated alpha blending weight $A_{i}=\sum_{p\in\mathcal{P}_{i}}a_{i}(p)$, where $a_{i}(p)$ is the blending weight of $g_{i}$ at pixel pp. The influence region is defined as a square centered at $(x_{i},y_{i})$ with half-extent $R_{i}=\left\lfloor\frac{\sqrt{C_{i}}}{2}\right\rfloor$, and the corresponding pixel set $\mathcal{P}_{i}$ includes all pixels within this region. We then approximate the average fused perceptual error for $g_{i}$ as
@@
SA [SplattingAvatar:CVPR2024] 27.48 0.9329 0.1046 24.93 0.8938 0.1478 26.14 0.9238 0.1145 26.02 0.8922 0.1821
MGA [chen2024monogaussianavatar] 27.34 0.9351 0.0859 27.91 0.9277 0.1113 26.52 0.9333 0.0900 27.04 0.9171 0.1225
GA [qian2023gaussianavatars] 26.98 0.9378 0.0851 24.62 0.9073 0.1333 25.74 0.9279 0.0965 25.08 0.8879 0.1763
FA [xiang2024flashavatar] 27.90 0.9357 0.0563 26.19 0.9055 0.0947 26.96 0.9331 0.0555 26.83 0.9110 0.0774
RGBA [li2025rgbavatar] 28.41 0.9493 0.0518 26.67 0.9217 0.0883 27.13 0.9438 0.0550 26.72 0.9277 0.0821
Fate [zhang2025fate] 28.33 0.9446 0.0508 28.36 0.9287 0.0776 27.12 0.9364 0.0558 27.18 0.9187 0.0849
@@ 3.4 Training Objective and Optimization
RGB Loss. We supervise the rendered images using a combination of $\mathcal{L}_{1}$ term and D-SSIM term following [kerbl3Dgaussians]. To better capture high-frequency facial details, we further introduce the perceptual loss $\mathcal{L}_{\mathrm{vgg}}$, which is activated only in the second half of training to balance performance and efficiency. The overall image loss is defined as:
@@
where $\lambda_{3}=0.01$ and $\lambda_{4}=0.001$. Following GaussianAvatars [qian2023gaussianavatars], we also employ a position loss and a scale loss to regularize the spatial alignment and local scale of the 3D Gaussians with respect to their parent triangles.
@@
Optimization. We employ Adam [kingma2014adam] for parameter optimization, using identical hyperparameter settings across all subjects. The learning rate is set to $1\times 10^{-4}$ for the UV-Adaptive Soft Binding framework, while the learning rates for the remaining parameters follow those used in 3D Gaussian Splatting. In addition to the Gaussian Splatting parameters, we also finetune the expression, translation and joint rotation parameters of FLAME for each timestep as done in GaussianAvatars. As each Gaussian is bound to a mesh triangle, we do not observe noticeable floating Gaussians and thus skip opacity resetting in ADC.

### v3

@@ Abstract
Reconstructing high-fidelity and animatable 3D head avatars from monocular videos remains a challenging yet essential task. Existing methods based on 3D Gaussian Splatting typically bind Gaussians to mesh triangles and model deformations solely via Linear Blend Skinning, which results in rigid motion and limited expressiveness. Moreover, they lack specialized strategies to handle frequently occluded regions (e.g., mouth interiors, eyelids). To address these limitations, we propose STAvatar, which consists of two key components: (1) a UV-Adaptive Soft Binding framework that leverages both image-based and geometric priors to learn per-Gaussian feature offsets within the UV space. This UV representation supports dynamic resampling, ensuring full compatibility with Adaptive Density Control (ADC) and enhanced adaptability to shape and textural variations. (2) a Temporal ADC strategy, which first clusters structurally similar frames to facilitate more targeted computation of the densification criterion. It further introduces a novel fused perceptual error as clone criterion to jointly capture geometric and textural discrepancies, encouraging densification in regions requiring finer details. Extensive experiments on four benchmark datasets demonstrate that STAvatar achieves state-of-the-art reconstruction performance, especially in capturing fine-grained details and reconstructing frequently occluded regions. Our project is available at https://jiankuozhao.github.io/STAvatar/.
@@ 1 Introduction
Reconstructing animatable and photo-realistic 3D head avatars has been a longstanding problem in computer vision and graphics [3, 22, 29, 34, 26, 28], driven by growing demands in AR/VR, telepresence, digital humans and interactive media. To achieve high-quality avatar reconstruction, conventional methods typically rely on expensive multi-camera systems [7, 32] which require complex manual setups and are impractical for consumer-level deployment. In contrast, monocular-based approaches [25, 39, 6] aim to produce high-fidelity results using a single consumer-grade camera, enabling broader accessibility in immersive and personalized virtual environments.
@@
Recent years have witnessed the emergence of 3D Gaussian Splatting (3DGS) [8] as a powerful volumetric representation, delivering both high-fidelity rendering and real-time performance in neural scene reconstruction. Building on its success in static scenes, many works have explored its application to animatable 3D head avatar reconstruction. These methods typically follow a two-stage pipeline: a parametric head-tracking algorithm [20, 44] is first employed to estimate the facial pose, expression, and shape for each frame; subsequently, Gaussian primitives are bound to mesh triangles and driven by a deformation field, which is typically implemented using Linear Blend Skinning (LBS) [19]. During training, each Gaussian’s parameters are optimized via differentiable rendering to minimize photometric loss. Despite achieving enhanced visual fidelity and animation realism, existing Gaussian-based approaches still face several open challenges. To address these limitations, we propose STAvatar, as shown in Fig. 1, and provide a detailed explanation below.
@@
First, existing animatable reconstruction methods often suffer from rigid connection between Gaussian primitives and mesh triangles, resulting in poor handling of non‑rigid and fine‑scale deformations. Fig. 2(a) illustrates this simple hard binding framework [35, 24, 33] in which each Gaussian deforms solely according to the LBS of its parent triangle. Although this enables basic animation control, it fails to model fine-grained deformations, as the Gaussians remain relatively static within the triangle’s local coordinate frame. This limitation stems from the gap between mesh and Gaussian representations. While meshes inherently encode topology and support native skinning for deformation, Gaussians are structureless and rely on externally defined deformation fields to capture complex motions and fine details such as facial wrinkles. Some methods [1, 30, 5] attempt to enhance each Gaussian’s deformation using a fixed-dimension multilayer perceptron (MLP). However, this design requires predefining a fixed number of Gaussians, which is suboptimal since different identities often demand varying Gaussian densities. Besides, it neglects spatial context and may be incompatible with the Adaptive Density Control (ADC) inherent to 3DGS. To enhance deformation ability while preserving the flexibility of ADC, STAvatar proposes a UV-Adaptive Soft Binding framework, which integrates the LBS for coarse deformation and a dual-branch network for fine-detail recovery. By generating a feature offset map in UV space, STAvatar leverages spatial context to model the Gaussian deformation field and enables arbitrary-resolution sampling to support ADC.
@@
Second, the vanilla ADC in 3DGS is tailored for static scene reconstruction and fails to handle transiently visible or frequently occluded regions in dynamic avatar reconstruction. As illustrated in Fig. 2(b), certain regions (e.g., mouth interiors) are only visible in a subset of frames and contribute minimally to rendering when occluded. This results in low average criterion-such as the positional gradient used in 3DGS-which impedes sufficient densification. In addition, as shown in Fig. 2(c), the positional gradient in 3DGS only captures geometric discrepancies while neglecting texture details [21], which leads to the omission of regions with high texture errors and the failure to add Gaussians to these regions. To address these limitations of ADC, STAvatar propose a Temporal Adaptive Density Control strategy incorporating FLAME-Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE-AP). As shown in Fig. 2(b), FTC partitions frames into structure‑coherent clusters to encourage frequently occluded regions remain visible within a cluster, facilitating targeted refinement. FPE‑AP replaces positional gradients by directly estimating each Gaussian’s fused perceptual error, which jointly considers both geometric and textural discrepancies. Besides, FPE-AP selects Gaussians exhibiting the highest instantaneous errors across iterations to better capture regions with large peak errors.
@@ 2.1 Gaussian Head Reconstruction
3DGS [8] has recently emerged as a flexible and efficient approach for 3D scene representation. By employing a large number of anisotropic Gaussian primitives—each parameterized by geometry and appearance attributes—it achieves high-quality reconstructions, making it particularly well-suited for detailed human head modeling. To enable avatar driving, typical methods [19, 16, 25, 23, 24, 35] adopt a hard-binding framework that associates Gaussians with mesh triangles to transfer mesh motion via LBS, which, however, fails to bridge the inherent gap between mesh and Gaussian representations, leading to suboptimal performance in handling complex motions and expressions. Alternatively, other approaches [30, 1, 31, 14] leverage MLP to predict per-Gaussian attribute offsets, thereby achieving a more flexible soft-binding framework. However, these methods often neglect ADC by using a fixed number of Gaussians across different identities, which limits expressiveness and reconstruction quality. In addition, several studies [17, 12, 43] propose to represent avatars with a base model and a set of expression blendshapes. Although these methods can achieve high training efficiency, they often suffer from low fidelity or produce blurred results under large pose variations, since the learned blendshapes may not adequately capture diverse pose-dependent deformations.
@@ 2.2 Adaptive Density Control of 3DGS
ADC strategy is a key module in 3DGS which dynamically clones, splits and prunes Gaussians to balance detail preservation and representation compactness. The vanilla ADC selects primitives for densification based on average positional gradients and Gaussian scales. However, this heuristic often results in underfitting and blurred details, as it neglects texture-related cues and only focus average criterion. To overcome these limitations, some recent works [21, 37, 10, 18, 41] revise the densification criterion to incorporate richer information, thereby achieving finer details in underrepresented regions. Besides, optimization-theoretic approaches have been proposed to establish more principled rules for cloning and splitting: SteepGS [27] leverages steepest-descent analysis to trigger Gaussian splits, whereas 3DGS-MCMC [9] formulates density control as a Markov Chain Monte Carlo (MCMC) sampling process with $L_{1}$-regularized pruning. More recently, Efficient Density Control [2] introduces long-axis splitting operations and recovery-aware pruning strategies to reduce redundant Gaussians and accelerate convergence, while Gradient-Direction-Aware ADC [42] incorporates gradient coherence and angular alignment to more effectively determine which Gaussians should be cloned or split. However, as these methods are primarily tailored for static scene reconstruction, they tend to under-fit transiently visible regions and fail to jointly consider both geometric and textural discrepancies.
@@ 3 Method
We first introduce the general Gaussian avatar reconstruction pipeline in Sec. 3.1. Next, we present a UV-Adaptive Soft Binding framework in Sec. 3.2. The Temporal Adaptive Density Control strategy, which integrates FLAME‑Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE‑AP), is discussed in Sec. 3.3. Finally, the detailed optimization process is described in Sec. 3.4.
@@ 3.1 Preliminaries
Initialization. Following [19], we initialize Gaussian primitives $g_{i}$ based on the tracked FLAME model. Each $g_{i}$ is bound to a mesh triangle and parameterized by a center position $\bm{\mu}$, a scale matrix $\bm{S}$, and a rotation matrix $\bm{R}$ defined on the canonical FLAME mesh [13]. The 3D Gaussian density is defined as:
@@ 3.3 Temporal Adaptive Density Control
To estimate each Gaussian’s 2D projection for fused perceptual error calculation, we record its screen-space center $(x_{i},y_{i})$, total number of covered pixels $C_{i}$, and accumulated alpha blending weight $A_{i}=\sum_{p\in\mathcal{P}_{i}}a_{i}(p)$, where $a_{i}(p)$ is the blending weight of $g_{i}$ at pixel $p$. The influence region is defined as a square centered at $(x_{i},y_{i})$ with half-extent $R_{i}=\left\lfloor\frac{\sqrt{C_{i}}}{2}\right\rfloor$, and the corresponding pixel set $\mathcal{P}_{i}$ includes all pixels within this region. We then approximate the average fused perceptual error for $g_{i}$ as
@@
SA [24] 27.48 0.9329 0.1046 24.93 0.8938 0.1478 26.14 0.9238 0.1145 26.02 0.8922 0.1821
MGA [1] 27.34 0.9351 0.0859 27.91 0.9277 0.1113 26.52 0.9333 0.0900 27.04 0.9171 0.1225
GA [19] 26.98 0.9378 0.0851 24.62 0.9073 0.1333 25.74 0.9279 0.0965 25.08 0.8879 0.1763
FA [30] 27.90 0.9357 0.0563 26.19 0.9055 0.0947 26.96 0.9331 0.0555 26.83 0.9110 0.0774
RGBA [12] 28.41 0.9493 0.0518 26.67 0.9217 0.0883 27.13 0.9438 0.0550 26.72 0.9277 0.0821
Fate [35] 28.33 0.9446 0.0508 28.36 0.9287 0.0776 27.12 0.9364 0.0558 27.18 0.9187 0.0849
@@ 3.4 Training Objective and Optimization
RGB Loss. We supervise the rendered images using a combination of $\mathcal{L}_{1}$ term and D-SSIM term following [8]. To better capture high-frequency facial details, we further introduce the perceptual loss $\mathcal{L}_{\mathrm{vgg}}$, which is activated only in the second half of training to balance performance and efficiency. The overall image loss is defined as:
@@
where $\lambda_{3}=0.01$ and $\lambda_{4}=0.001$. Following GaussianAvatars [19], we also employ a position loss and a scale loss to regularize the spatial alignment and local scale of the 3D Gaussians with respect to their parent triangles.
@@
Optimization. We employ Adam [11] for parameter optimization, using identical hyperparameter settings across all subjects. The learning rate is set to $1\times 10^{-4}$ for the UV-Adaptive Soft Binding framework, while the learning rates for the remaining parameters follow those used in 3D Gaussian Splatting. In addition to the Gaussian Splatting parameters, we also finetune the expression, translation and joint rotation parameters of FLAME for each timestep as done in GaussianAvatars. As each Gaussian is bound to a mesh triangle, we do not observe noticeable floating Gaussians and thus skip opacity resetting in ADC.
@@ 4.1 Experimental Setup
Implementation Details. We conduct both training and inference on a single NVIDIA RTX 3090 GPU. Each identity’s video is trained for 6 epochs. All videos are cropped and resized to a resolution of $512\times 512$. A robust matting method [15] is applied to extract the foreground with the background set to white. For avatar initialization, we adopt VHAP [20] for monocular head tracking.
Dataset. We evaluate our method on four datasets, including INSTA [45], PointAvatar [40], Nerface [4], and HDTF [38], comprising a total of 22 identities. For PointAvatar, we follow the original data split strategy in [40], while for the other datasets, the last 450 frames are reserved as the test set.
Baselines. We compare our method against six state-of-the-art Gaussian avatar reconstruction approaches: GaussianAvatars (GA) [19], FateAvatar (Fate) [35], RGAvatar (RGBA) [12], FlashAvatar (FA) [30], MonoGaussianAvatar (MGA) [1], and SplattingAvatar (SA) [24]. Note that simply equalizing or increasing training epochs across methods may not yield optimal results. Accordingly, we follow the settings in the original papers, setting 10 epochs for GA, Fate, and RGBA, 20 for FA, 30 for SA, and 100 for MGA to allow each method to converge properly. Since GA is originally designed for multi-view reconstruction, we adopt various settings from original papers for fair comparison.
4.2 Evaluation
Quantitative Results. We select PSNR, SSIM, and LPIPS [36] as evaluation metrics to comprehensively assess the visual quality of the reconstructed avatars. As shown in Tab. 1, our method consistently outperforms existing state-of-the-art approaches across most datasets and metrics, achieving notable improvements in reconstruction quality. In particular, our method attains the highest SSIM scores and the lowest LPIPS values on all four datasets, demonstrating its strong ability to preserve both geometric accuracy and perceptual fidelity.